OpenStudy (anonymous):

How do I solve and justify 3x-1=12x-4

OpenStudy (anonymous):

solve means find the unknown so find x by isolating it in the equation. do this by combining all terms with x using algebra. 3x-1=12x-4 (take 3x to same side as 12x) -1=12x-3x-4 (take -4 to other side and combine 3x and 12x) -1+4=9x 3=9x (divide both sides of equation by 9 to isolate x) x= 1/3 to justify, plug x=1/3 back into original equation and see if left side equals right side (if it does, you did the question right) 3x-1=12x-4 3(1/3)-1=12(1/3)-4 3/3 - 1 = 12/3 - 4 1-1=4-4 0=0 (justified since both sides are equal)

OpenStudy (anonymous):

Thank youu can you help me with this one 5n=-4n+18

OpenStudy (anonymous):

yeah sure

OpenStudy (anonymous):

you essentially do the same thing... 5n=-4n + 18 (bring n terms to same side) 5n + 4n = 18 9n = 18 (divide both sides by 9) n=18/9 n=2 to verify, plug n back into equation... 5n=-4n+18 5(2) = -4(2)+18 10 = -8+18 10=10 (both sides are equal so x really is 2)